# Pangu DeepDiver: Adaptive Search Intensity Scaling via Open-Web Reinforcement Learning

### v1

@@ 3.2 DeepDiver
Building upon the WebPuzzle, we showcase its efficacy within an RL framework designed to explore the information-seeking behavior of LLMs. In this section, we present our method, DeepDiver. DeepDiver ultilize the procedure of cold-start supervised fine-tuning (SFT) followed by reinforcement learning (RL), while incorporates a carefully designed reward assignment and scheduling mechanism to maintain stable RL training.
Initialization of Reasoning and Searching To equip DeepDiver with essential reasoning and searching capabilities for WebPuzzle, we implement a cold-start supervised fine-tuning process using diverse data: 2,000 WebPuzzle samples across difficulty levels, 300 real-user questions from our deployed smart assistant, 2,200 general reasoning problems from recent studies (Jiang et al., 2024; Min et al., 2024; Xu et al., 2023; Yang, 2023; Zhang et al., 2023), and 1,000 real-user queries concatenated with retrieved documents. This dataset distills responses from DeepSeek-R1, establishing DeepDiver’s foundational abilities to iteratively search and reason over retrieved documents. The distillation prompt configuration is detailed in Appendix C.3.
GRPO With Iterative RAG After SFT, we enhance DeepDiver by extending GRPO (Shao et al., 2024) with iterative RAG. As shown in Figure LABEL:fig:webdiver, the model iteratively performs reasoning and searching until reaching an acceptable answer, following the pipeline in Section 2.1. We apply a loss mask to distinguish model-generated from externally retrieved tokens, with GRPO updating parameters based solely on model-generated content.
Extra Search Call Rewards Beyond standard format and accuracy rewards in GRPO, we introduce an extra reward to encourage search engine use for complex problems. When no search-free rollouts solve a problem but at least one search-enabled rollout succeeds, we assign an additional reward of $1.0$ to the successful search-enabled solutions. This ensures the model learns to leverage external tools when necessary. The formal definition appears in Appendix C.2.
Loose and Strict Rewards Our reward function employs LLM-based graders in a two-stage training approach that transitions from loose to strict grading: the loose grader assigns scores from 1 to 10 (scores $\geq 6$ yield 1.0 reward), particularly benefiting early training as shown in Section 5.3. The strict grader conducts three evaluation rounds, requiring at least 2 of 3 positive judgments. Both grader definitions appear in Appendix C.1.
@@ 4.1 Setup
Data Mixture and Selection
Due to computational constraints and the capability limits of the 7B model, we do not train DeepDiver on the full WebPuzzle dataset. Instead, we carefully select and mix 7k samples for training, which are evenly split into 2k and 5k subsets based on tagging categories. The 2k subset is used as part of the cold-start SFT data, as described in Section 3, while the remaining 5k samples are reserved for RL training. Empirical results show that this mixture strategy strikes a balance between computational efficiency and model effectiveness. Detailed statistics of the mixture data are presented in Table 8.
Benchmark Datasets
We evaluate the performance of all models using the following closed-ended benchmarks:
  • C-SimpleQA-500: A randomly sampled subset of 500 instances from C-SimpleQA (Wei et al., 2024), C-SimpleQA is a Chinese-translated version designed to assess the factuality of language models. While not explicitly designed for complex or real-time question answering, we utilize C-SimpleQA to explore the impact of web search scaling on simpler information-seeking tasks.
  • FRAMES-zh-230: A subset of the FRAMES (Krishna et al., 2024) benchmark with 230 samples that requires multi-hop information-seeking. The queries are translated into Chinese, and annotators verify whether the golden answer can be retrieved via interactions with our web search API. Only test cases where the golden answer is reachable are included in the evaluation.
  • Bamboogle-zh-71: A subset of the Bamboogle (Press et al., 2023) benchmark consists of 71 samples, processed similarly to the FRAMES-zh-230 dataset.
Baseline Models
For all trainable baselines, we use Qwen2.5-7B-Instruct (Team, 2024) and Pangu-7B-Reasoner (Tang et al., 2024) as the backbone model. The training-free baseline adopt the QwQ-32B (Team, 2025), GPT-4o (OpenAI, 2024b) and DeepSeek-R1 (DeepSeek-AI et al., 2025). The following methods are evaluated:
  • Prompted without Web Search: In this setup, LLMs respond to problems based on a single round of prompting without web access. The model generates answers based solely on its pre-existing knowledge. We include off-the-shelf LLMs such as Qwen2.5-7B-Instruct (Team, 2024; Yang et al., 2024), QwQ-32B (Team, 2025), GPT-4o (OpenAI, 2024b), DeepSeek-R1 (DeepSeek-AI et al., 2025) and Pangu-7B-Reasoner (Tang et al., 2024).
  • Prompted with Iterative RAG: In this approach, LLMs answer problems through multiple rounds of retrieval and reasoning, using a real-world open-web search engine. The baselines from the “Prompted without Web Search Methods” are tested using the same iterative RAG workflow in our approach (see Appendix C.3 for implementation details).
Evaluation Metrics
@@
Qwen2.5-7B-Ins. (Team, 2024) 7.4 28.4 14.1 19.7
Pangu-7B-Reasoner (Tang et al., 2024) 15.0 36.3 20.4 27.2
GPT-4o (OpenAI, 2024b) 14.2 61.8 51.7 52.6
QwQ-32B (Team, 2025) 21.9 51.3 36.5 54.5
DeepSeek-R1 (DeepSeek-AI et al., 2025) 32.7 74.6 63.8 73.2
@@
DeepDiver-Pangu-7B (Ours) (2.89) (2.61) (3.05) (2.72)
Table 2: Numbers in () indicate average search call rounds per example, with each round potentially involving 1$\sim$5 search queries. Results show average accuracy across 3 runs.
We employ the strict grader outlined in Section 3.2 as our evaluation metric. Unlike conventional LLM-as-a-judge (Zheng et al., 2023) methods, which directly assess the quality of the generated response, our grader considers both the reference answer and the provided checklists, resulting in a more robust evaluation. Specifically, the strict grader evaluates the generated response by comparing it to the reference answer and the checklists over three rounds. In each round, the evaluator determines whether the model’s output matches the reference answer and aligns with the requirements specified in the checklists. If the evaluator deems the response correct in at least two out of the three rounds, the result is classified as correct; otherwise, it is considered incorrect. The accuracy rate, derived from this evaluation process, is reported as the primary metric in our results.
@@ 4.2 How does DeepDiver’s Performance Compare to Baselines?
DeepDiver-Pangu-7B shows similar improvements. While R1-distilled Pangu-7B quickly hits performance bottlenecks (dropping 5.7 points on Bamboggle compared to the cold-start model), DeepDiver-Pangu-7B breaks through these limitations, showing substantial improvements across all benchmarks and achieving exceptional performance (38.1) on WebPuzzle. In conclusion, both Pangu-7B-reasoner and Qwen DeepDiver demonstrate competitive performance against high-performing models like DeepSeek-R1 and QwQ. This highlights DeepDiver’s capability to effectively search for and reason over relevant information and solve complex reasoning tasks through search intensity scaling.
@@ 4.3 What is the Relationship between Search Intensity and Performance?
Search intensity is strongly correlated with performance improvements, as increases in search frequency and depth during the RL phase consistently lead to better outcomes. Figure 4 illustrates this relationship during the training phase, showing a clear trend: as search engine calls increase, so do training rewards. For testing results, despite SFT’s progress compared with the prompting methods, the model faces a performance bottleneck to adapt to more challenging problems, still lagging behind off-the-shelf APIs with a large margin. In contrast, our RL-based DeepDiver-Qwen2.5-7B promotes higher search intensity with an average of 2.51 search and reasoning rounds, substantially higher than the SFT model’s 1.75. Similar gains appear in DeepDiver-Pangu-7B, where increased search rounds ($1.84\rightarrow 2.89$) correspond to performance improvements ($30.3\rightarrow 38.1$). This searching intensity scaling enables models to explore and verify more relevant information, enhancing their ability to tackle complex problems.
@@ 4.4 Can DeepDiver Generalize from Open-web Training to OOD Wiki-based Problem?
Training with WebPuzzle, DeepDiver demonstrates strong generalization capabilities and performance improvements on Wiki-based problems. DeepDiver shows impressive generalization on Wiki-based benchmarks despite not being specifically trained for these tasks. Both DeepDiver-Qwen2.5-7B and DeepDiver-Pangu-7B significantly outperform their distilled variants and demonstrate substantial improvements over cold-start models. While DeepSeek-R1 performs well on Wiki-based problems without web search, it shows modest gains when combined with the iterative RAG pipeline. This suggests DeepSeek-R1 may have already internalized the necessary knowledge for Wiki-based problems, highlighting the importance of our proposed WebPuzzle benchmark. We further investigate this hypothesis through isolated tests on information seeking and verification in Section 5.1.
@@ 5 Analysis
This section focuses on the Qwen2.5-7B-Instruct model, a simpler LLM that lacks advanced reasoning capabilities compared to the Pangu-Reasoner, to evaluate the effectiveness of WebPuzzle and DeepDiver. We analyze several key aspects, including isolated evaluations of information-seeking behavior, comparisons with concurrent related work, the design of the reward function, and the model’s generalization to open-ended problems. We also demonstrate additional analyses—such as the relationship between search intensity and problem difficulty, statistics of information-seeking behavior across different training and testing environments, comparisons between human and DeepDiver performance, and detailed case studies.
@@ 5.1 Isolation Testing of Information-Seeking
Figure 5: Pairwise comparisons between DeepDiver 7B and baselines. We filter out problems solvable by both models using internal knowledge alone, then report accuracy on the remaining problems. DeepDiver exhibits exceptional information-seeking and adaptive search intensity scaling capabilities.
@@
Results DeepDiver exhibits exceptional information-seeking capabilities, comparable to all baselines on problems that cannot be solved by internal knowledge alone. While our 7B DeepDiver initially trails behind 671B baselines in full-set tests, results shift when isolating information-seeking behavior. As Figure 5 shows, on problems challenging even for larger models, DeepDiver demonstrates competitive performance across all benchmarks. Notably, it outperforms DeepSeek-R1 across all domains, with a 5.1-point lead on WebPuzzle. This suggests our 7B model’s limitations in full-dataset performance stem primarily from its limited internal knowledge. However, when tackling problems requiring external information search and verification, DeepDiver’s information-seeking capability demonstrates strength in addressing real-world open-web problems.
@@ 5.2 Comparisons with Wiki-based Methods
To highlight limitations of wiki-based training queries and search engines, we compare DeepDiver with prior wiki-based methods. Despite being trained entirely in Chinese, we evaluate DeepDiver on English benchmarks with English search engines to demonstrate its robustness and generalizability.
@@
R1-Searcher (Song et al., 2025) 13.7 (1.9) 46.7 (2.0) 25.3 (1.9) 57.9 (2.3)
DeepResearcher (Zheng et al., 2025) 15.0 (7.5) 53.9 (7.1) 33.6 (7.2) 56.6 (4.4)
DeepDiver (ours) 26.1 (14.7) 56.8 (9.1) 32.0 (14.2) 58.4 (10.4)
@@
Setup We use R1-Searcher (Song et al., 2025) and DeepResearcher (Zheng et al., 2025) as baselines—both trained in English using Wiki-based corpora. Search engine settings appear in Appendix C.5. For evaluation, we translate WebPuzzle into English via Qwen Max (Team, 2024), use the full Bamboogle dataset (Press et al., 2023) (125 examples), and randomly sample 300 examples from FRAMES (Krishna et al., 2024) and HotpotQA (Yang et al., 2018). For fairness, we report accuracy based on the average judgment across all methods.
Results Despite the language gap, DeepDiver–trained in a real-world Chinese internet setting using WebPuzzle queries–outperforms Wiki-based baselines on most tasks, underscoring the strength of SIS. As shown in Table 3, DeepDiver significantly outperforms DeepResearchers on WebPuzzle-en with an 11.1-point lead, while maintaining strong results on Wiki-based datasets despite no English training for information-seeking. We attribute this to DeepDiver’s use of SIS, which enables intensive information retrieval and verification rather than relying on limited internal knowledge or language-specific constraints. R1-Searcher and DeepResearcher make significantly fewer search calls than DeepDiver due to their ”cleaner” and more constrained training environments, leading to poorer real-world performance when facing the noise and complexity of open information-seeking tasks. For additional results, including individual judge assessments, see Table 9.
@@ 5.3 Tolerance of the Reward Function
During DeepDiver’s RL training, we observed a reward plateau after approximately 80 optimization steps. We investigated potential factors including learning rate scheduler, exploration diversity, environmental instability, and gradient issues, but found no obvious problems. We therefore focused on the reward function design as a potential cause of the performance plateau.

### v2

@@ 3.2 DeepDiver
Building upon the WebPuzzle, we showcase its efficacy within a RL framework designed to explore the information-seeking behavior of LLMs. In this section, we present our method, DeepDiver. DeepDiver ultilize the procedure of cold-start supervised fine-tuning (SFT) followed by reinforcement learning (RL), while incorporates a carefully designed reward assignment and scheduling mechanism to maintain stable RL training.
Figure 3: DeepDiver overview. (a) Rollout Generation: DeepDiver iteratively reasons, retrieves evidence, and answers WebPuzzle queries, then receives rewards based on comparison with ground truth. (b) RL Updates: Retrieved text is masked during loss calculation, and the LLM is refined via GRPO using advantages $A_{i}$ derived from rewards $r_{i}$.
Initialization of Reasoning and Searching To equip DeepDiver with essential reasoning and searching capabilities for WebPuzzle, we implement a cold-start supervised fine-tuning process using diverse data: 2,000 WebPuzzle samples across difficulty levels, 300 real-user questions from our deployed smart assistant, 2,200 general reasoning problems from recent studies [Slow_Thinking_with_LLMs_1, Slow_Thinking_with_LLMs_2, xu2023cvalues, Firefly, zhang2023chinese], and 1,000 real-user queries concatenated with retrieved documents. This dataset distills responses from DeepSeek-R1, establishing DeepDiver’s foundational abilities to iteratively search and reasoning over retrieved documents. The distillation prompt configuration is detailed in Appendix E.6.
GRPO With Iterative RAG After SFT, we enhance DeepDiver by extending GRPO [shao2024deepseekmathpushinglimitsmathematical] with iterative RAG. As shown in Figure 3, the model iteratively performs reasoning and searching until reaching an acceptable answer, following the pipeline in Section 2.1. We apply a loss mask to distinguish model-generated from externally retrieved tokens, with GRPO updating parameters based solely on model-generated content.
Extra Search Call Rewards Beyond standard format and accuracy rewards in GRPO, we introduce an extra reward to encourage search engine use for complex problems. When no search-free rollouts solve a problem but at least one search-enabled rollout succeeds, we assign an additional reward of $1.0$ to the successful search-enabled solutions. This ensures the model learns to leverage external tools when necessary. The formal definition appears in Appendix E.3.
Loose and Strict Rewards Our reward function employs LLM-based graders in a two-stage training approach that transitions from loose to strict grading: the loose grader assigns scores from 1 to 10 (scores $\geq 6$ yield 1.0 reward), particularly benefiting early training as shown in Section 5.4. The strict grader conducts three evaluation rounds, requiring at least 2 of 3 positive judgments. Both grader definitions appear in Appendix E.2.
@@ 4.1 Setup
Data Mixture and Selection Due to computational constraints and capability limits of the 7B model, we train DeepDiver on a carefully selected mixture of 7k WebPuzzle samples rather than the full dataset. We evenly split these into 2k samples for cold-start SFT (Section 3) and 5k for RL training. This mixture strategy balances computational efficiency and model effectiveness. Detailed statistics appear in Table 8.
Benchmark Datasets and Baseline Models We evaluate performance using closed-ended Chinese benchmarks including C-simpleQA-500 [wei2024measuringshortformfactualitylarge, he2024chinesesimpleqachinesefactuality], FRAMES-zh-230 [krishna2024factfetchreasonunified], BamBoogle-zh-71 [press2023measuringnarrowingcompositionalitygap], and our proposed WebPuzzle (detailed in Appendix E.4). For trainable baselines, we use Qwen2.5-7B-Instruct [qwen2.5] and Pangu-7B-Reasoner [tang2024rethinking] as backbone models. Training-free baselines include QwQ-32B [qwq32b], GPT-4o [openai_gpt4o] and DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability]. We evaluate methods including Prompted without Web Search, Prompted with Iterative RAG, and R1-Distillation (detailed in Appendix E.5). Our evaluation uses the strict grader from Section 3.2, which considers both reference answers and checklists for more robust assessment than conventional LLM-as-a-judge [zheng2023judgingllmasajudgemtbenchchatbot] approaches. Grader details appear in Appendix E.2.
@@
Qwen2.5-7B-Ins. [qwen2.5] 7.4 28.4 14.1 19.7
Pangu-7B-Reasoner [tang2024rethinking] 15.0 36.3 20.4 27.2
GPT-4o [openai_gpt4o] 14.2 61.8 51.7 52.6
QwQ-32B [qwq32b] 21.9 51.3 36.5 54.5
DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] 32.7 74.6 63.8 73.2
@@
DeepDiver-Pangu-7B (2.89) (2.61) (3.05) (2.72)
Table 1: Numbers in () indicate average search call rounds per example, with each round potentially involving 15 search queries. Results show average accuracy across 3 runs.
@@ 4.2 How does DeepDiver’s Performance Compare to Baselines?
DeepDiver-Pangu-7B shows similar improvements. While R1-distilled Pangu-7B quickly hits performance bottlenecks (dropping 5.7 points on Bamboggle compared to the cold-start model), DeepDiver-Pangu-7B breaks through these limitations, showing substantial improvements across all benchmarks and achieving exceptional performance (38.1) on WebPuzzle. In conclusion, both Pangu-7B-reasoner and Qwen DeepDiver demonstrate competitive performance against high-performing models like DeepSeek-R1 and QwQ. This highlights DeepDiver’s capability to effectively search for and reason over relevant information, and solve complex reasoning tasks through search intensity scaling.
@@ 4.3 What is the Relationship between Search Intensity and Performance?
Search intensity is strongly correlated with performance improvements, as increases in search frequency and depth during the RL phase consistently lead to better outcomes. Figure 4 illustrates this relationship during the training phase, showing a clear trend: as search engine calls increase, so do training rewards. For testing results, despite SFT’s progress compared with the prompting methods, the model faces a performance bottleneck to adapt to more challenging problems, still lagging behind off-the-shelf APIs with large margin. In contrast, our RL-based DeepDiver-Qwen2.5-7B promotes higher search intensity with an average of 2.51 search and reasoning rounds, substantially higher than the SFT model’s 1.75. Similar gains appear in DeepDiver-Pangu-7B, where increased search rounds ($1.84\rightarrow 2.89$) correspond to performance improvements ($30.3\rightarrow 38.1$). This searching intensity scaling enables models to explore and verify more relevant information, enhancing their ability to tackle complex problems.
@@ 4.4 Can DeepDiver Generalize from Open-web Training to OOD Wiki-based Problem?
Training with WebPuzzle, DeepDiver demonstrates strong generalization capabilities and performance improvements on Wiki-based problems. DeepDiver shows impressive generalization on Wiki-based benchmarks despite not being specifically trained for these tasks. Both DeepDiver-Qwen2.5-7B and DeepDiver-Pangu-7B significantly outperform their distilled variants and demonstrate substantial improvements over cold-start models. While DeepSeek-R1 performs well on Wiki-based problems without web search, it shows modest gains when combined with iterative RAG pipeline. This suggests DeepSeek-R1 may have already internalized the necessary knowledge for Wiki-based problems, highlighting the importance of our proposed WebPuzzle benchmark. We further investigate this hypothesis through isolated tests on information seeking and verification in section 5.1.
@@ 5 Analysis
This section focuses on the Qwen2.5-7B-Instruct model, a simpler model comparing with the Pangu-7B-Reasoner. We analyze several key aspects, including isolated evaluations of information-seeking behavior, comparisons with concurrent related work, the design of the reward function, and the model’s generalization to open-ended problems. Additional analyses—such as the relationship between search intensity and problem difficulty, statistics of information-seeking behavior across different training and testing environments, comparisons between human and DeepDiver performance, and detailed case studies—are provided in Appendix A.
@@ 5.1 Isolation Testing of Information-Seeking
Figure 5: The comparison after removing cases answered correctly through internal knowledge.
@@
Results DeepDiver exhibits exceptional information-seeking capabilities, comparable to all baselines on problems that cannot be solved by internal knowledge alone. While our 7B DeepDiver initially trails behind 671B baselines in full-set tests, results shift when isolating information-seeking behavior. As Figure 5 shows, on problems challenging even for larger models, DeepDiver demonstrates competitive performance across all benchmarks. Notably, it outperforms DeepSeek-R1 across all domains, with a 5.1-point lead on WebPuzzle. This suggests our 7B model’s limitations in full-dataset performance stem primarily from its smaller size limiting internal knowledge. However, when tackling problems requiring external information search and verification, DeepDiver’s information-seeking capability demonstrates strength in addressing real-world open-web problems.
@@ 5.2 Comparisons with Wiki-based Methods
To highlight wiki-based training environments’ limitations, we compare DeepDiver with prior wiki-based methods. Despite being trained entirely in Chinese, we evaluate DeepDiver on English benchmarks with English search engines to demonstrate its robustness and generalizability.
@@
R1-Searcher [song2025r1searcherincentivizingsearchcapability] 13.7 (1.9) 46.7 (2.0) 25.3 (1.9) 57.9 (2.3)
DeepResearcher [zheng2025deepresearcher] 15.0 (7.5) 53.9 (7.1) 33.6 (7.2) 56.6 (4.4)
DeepDiver-Qwen 26.1 (14.7) 56.8 (9.1) 32.0 (14.2) 58.4 (10.4)
@@
Setup We use R1-Searcher [song2025r1searcherincentivizingsearchcapability] and DeepResearcher [zheng2025deepresearcher] as baselines—both trained in English using Wiki-based corpora. Search engine settings appear in Appendix E.8. For evaluation, we translate WebPuzzle into English via Qwen Max [qwen2.5], use the full Bamboogle dataset [press2023measuringnarrowingcompositionalitygap] (125 examples), and randomly sample 300 examples from FRAMES [krishna2024factfetchreasonunified] and HotpotQA [yang2018hotpotqadatasetdiverseexplainable]. For fairness, we report accuracy based on the average judgment across all methods.
Results Despite the language gap, DeepDiver–trained in a real-world Chinese internet setting using WebPuzzle queries–outperforms Wiki-based baselines on most tasks, underscoring the strength of SIS. As shown in Table 2, DeepDiver significantly outperforms DeepResearchers on WebPuzzle-en with 11.1 point leads, while maintaining strong results on Wiki-based datasets despite no English training for information-seeking. We attribute this to DeepDiver’s use of SIS, which enables intensive information retrieval and verification rather than relying on limited internal knowledge or language-specific constraints. R1-Searcher and DeepResearcher make significantly fewer search calls than DeepDiver due to their "cleaner" and more constrained training environments, leading to poorer real-world performance when facing the noise and complexity of open information-seeking tasks. For additional results, including individual judge assessments, see Table 9.
5.3 Emergence of the SIS
A natural concern is that SIS could be an artifact of reward shaping rather than a genuine behavior that emerges from training in a real web environment. To address this concern, we analyze whether the extra search-call reward introduced in Section E.3 consistently encourages the model to prefer search over no-search when both solve the task.
Setup
Theoretically, recall the bonus is only awarded when at least one search-enabled rollout succeeds and no search-free rollout succeeds for the same prompt group, i.e., it should not reward search when both search and no-search solve the task. Consequently, during training, we tracked the frequency of the extra bonus. Every 10 steps (448 trajectories per step), we counted how often the bonus fired and compared this against the evolving search intensity (average number of tool-use rounds per query).
Figure 6: The count and percentage of trajectories triggering the special auxiliary search reward (value=3.0) across RL training step ranges. The trigger frequency drops sharply (4.5% in 0–9 steps to 0.1% in 70–80 steps) and fades after step 30, confirming the reward’s transient role and supporting SIS as an emergent behavior.
Results
SIS is not merely reward-shaped but an emergent behavior developed during RL training, our reward design serves as a transient early scaffold (not persistent incentive) and differs fundamentally from existing works that explicitly encourages the tool calls. As shown in Figure 6, the special search reward exhibits a clear phasing-out trend: its trigger percentage drops from 4.5% (198 occurrences) in steps 0–9 to 0.1% (6 occurrences) in steps 70–80. Two key observations confirm its transient role: (1) Even in the earliest training phase, only a small fraction of trajectories (4.5%) received the reward, ruling out "over-rewarding"; (2) After step 30, the trigger percentage remains 1.1%, indicating the reward is essentially inactive in mid-to-late training. More importantly, Figure 4 shows that tool-use rounds grow sharply during steps 80–120—well after the auxiliary reward fades. This confirms SIS is not a reward-driven "phenomenon" but an emergent behavior: the model proactively leverages external tools to compensate for internal knowledge limitations, even without direct incentives.
@@ 5.4 Tolerance of the Reward Function
During DeepDiver-Qwen7B’s RL training, we observed a reward plateau after approximately 80 optimization steps. We investigated potential factors including learning rate scheduler, exploration diversity, environmental instability, and gradient issues, but found no obvious problems. We therefore focused on the reward function design as a potential cause of the performance plateau.